1. In lecture, we derived the electric field a height z above the center of a thin ring of charge with constant charge per unit length λ. Lets assumie here that λ > 0. Suppose a negative point charge q with mass m s placed a very small distance above the center of the ring. Show that the point charge undergoes simple harmonic motion and find the frequency of small oscillations. Hint: show that near the center of the ring the force on the point charge resembles Hookes law.
Answer #1

Electric field due to a charged ring at a distance 'z' from the center of the ring is given by

                                                         4περ(a +22)3/2

Where, Q (=2 \pi a \lambda) is total charge on the ring, a is radius of the ring and

                                                           -= 9x 109 Nm2 /c2 4πέο

Given, a negative charge -q is placed at a very small distance from the center of the ring. i.e. z<<a .

Therefore, we can neglect 'z' in the denominator,

                                                       E = AEC

So, the force on the negative charge is

                                                        F = -qΕ =- Qq 4πέρα3

                                                 or, m \, \frac{d^{2}z}{dt^{2}}= -\left ( \frac{Qq}{4 \pi \epsilon_{0} a^{3} } \right ) z

                                                 or, df2 = -(k/m)                     .........................................(1)

Where,                               * = Qq 4περα, , m is mass of the particle.

Equation 1 is the equation of SHM (Simple harmonic Motion).

Hence, Hence, the particle will do SHM with frequency,

                                           ( *) – u/41 =

For any doubt please comment.